Determine the first two terms of the following sequence: 8, 15, 24, 35


1
Expert's answer
2021-07-06T07:25:52-0400

15-8=7

24-15=9

35-24=11

So, difference sequence: 3, 5, 7, 9, 11

8-5=3

3-3=0

So, the original sequence: 0, 3, 8, 15, 24, 35